# Causal Interventions in Bond Multi-Dealer-to-Client Platforms

### v1

@@ 6 Generative versus discriminative models for causal interventions: empirical results
We evaluate two discriminative models: a logistic regression and a tree-based model, the Light Gradient Boosting Machine (LightGBM). Logistic regression serves as a simple benchmark that inherently respects monotonicity constraints on the predicted probability with respect to input features—particularly the quoted spread. In contrast, LightGBM is a powerful, widely used model known for its strong out-of-the-box predictive performance. Although it does not enforce monotonicity, we employ it to explore the predictive frontier of the problem, assessing the potential gains in accuracy when the monotonicity constraint is relaxed. A promising alternative involves neural network architectures that explicitly incorporate monotonicity constraints; see [25] for a related application in the context of derivatives pricing. These models aim to combine the principled structure of monotonic estimators with the flexibility and representational power of neural networks. Results from this line of research will be presented in future work [26].
@@
Finally, we include features specific to the request-for-quote (RfQ) level, corresponding to the variable $RF$ in the graphical model:
@@
We use a proprietary dataset of RfQs from BBVA, consisting of 102,437 RfQs, of which 5,738 are labeled as hits. The dataset is split into training, validation, and test sets, with the test set reserved for final model evaluation. The validation set is used for model selection and hyperparameter tuning in the discriminative ML models. For the logistic regression model, we tune the inverse regularization strength, setting it to 100 with an L2 penalty. For the LightGBM model, the best-performing hyperparameters identified are: column sample by tree = 1, learning rate = 0.01, minimum child samples = 50, number of estimators = 500, number of leaves = 15, and subsample = 0.6. The generative model does not require hyperparameter tuning and is trained using Maximum Likelihood Estimation (MLE) on the training set. Given the class imbalance in the dataset, we apply class weighting during training for both the generative and discriminative models. In addition, standard outlier cleaning techniques are applied to preprocess the dataset.
@@
A worst-case Balanced Brier Score (BBS) is $0.5$, while a perfect model achieves a score of 00. In terms of skill, the Balanced Brier Skill Score (BBSS) equals 00 when the model’s performance matches that of the benchmark, and reaches $1$ for a perfect model. This metric provides a succinct summary of the information conveyed by calibration plots, which group predicted probabilities into bins and compare the average predicted probability within each bin to the observed frequency of positive outcomes. Such plots—and, by extension, the Brier score—are particularly informative in settings where the predicted probabilities themselves, rather than just binary class labels, directly influence decision-making, as is the case in the business problems analyzed in this work.
The results for ROC-AUC and BBSS in the test set for the three models considered as well as the majority class benchmark are shown in Table 1.
Model ROC-AUC BBSS
@@
Table 1: Predictive scores: AUC-ROC and BBSS for the three models analyzed. We also show the metrics for the benchmark model, the majority class. The generative model achieves the best overall scores in both metrics.
The comparative analysis of the three models reveals important insights into both predictive performance and the quality of probability estimates. While LightGBM achieves the highest ROC-AUC at 0.743, its advantage over the generative model (0.742) is marginal and not practically significant. What makes the generative model particularly noteworthy is its ability to produce well-calibrated probabilities, as evidenced by its superior Balanced Brier Skill Score (BBSS) and the shape of its calibration curve, see Figure 3. Although none of the models achieve perfect calibration, the generative model clearly demonstrates the most consistent and reliable probabilistic output. This performance is particularly impressive given the relatively simple distributional assumptions underlying the model, suggesting that a generative approach can still serve effectively as a classifier while providing meaningful probability estimates.
@@
Fig 3: Calibration plots for the three evaluated models and the majority class baseline. The generative model exhibits the best calibration, as confirmed by its superior Balanced Brier Skill Score (BBSS) in Table 1
The relatively poor calibration of LightGBM is expected, given that it is not probabilistic by design and relies on post-hoc methods to estimate probabilities. Nonetheless, it is interesting that a state-of-the-art predictive model like LightGBM is unable to clearly outperform the generative model in either ROC-AUC or BBSS, a result that is also reinforced by inspecting the ROC curve in Figure 4. This outcome points to the strength of incorporating domain-specific structure into the modeling process. In this case, the generative model benefits from an explicit representation of the internal RfQ negotiation mechanics, such as the simultaneous bidding of multiple dealers and the client’s choice of the most competitive quote. Moreover, it has the capacity to integrate post-trade signals—such as the cover price or knowledge of being the second-best bid—which enhances its alignment with real-world market behavior.
@@
In contrast, logistic regression demonstrates the weakest overall performance, with an AUC of 0.684. However, one of its notable strengths—shared with the generative approach—is its adherence to the economic constraint of monotonicity with respect to the quoted spread, a property that holds by construction. While the linear form of logistic regression limits its capacity to capture the complexity and competitiveness inherent in the RFQ process, it remains widely used in the industry due to its simplicity and interpretability. By comparison, LightGBM is a more flexible and powerful predictive model but does not enforce the crucial monotonic relationship between spread and hit probability. This limitation becomes particularly problematic in data-sparse regions, such as for small or highly negative spreads. As illustrated in Figure 5, LightGBM exhibits erratic behavior in these areas, failing to reflect the expected increase in win probability as the quoted spreads become more attractive.
Overall, these findings underscore the importance of employing modeling strategies that incorporate the inherent structure and logic of the underlying process. While enhancements to purely predictive models—such as leveraging advanced architectures, fine-tuning hyperparameters, or enriching features with post-trade information—can yield incremental gains, the observed performance ceiling suggests that predictive accuracy alone is insufficient. In contrast, approaches that embed economic reasoning and structural constraints can deliver more robust and interpretable outcomes, without sacrificing predictive power, especially in complex financial settings such as negotiations within multi-dealer-to-client platforms.
One potential limitation of structurally informed models, however, is their inference time: generative approaches, in particular, tend to be slower than the other models evaluated. This drawback might be mitigated by the fact that low-latency inference is generally more critical in highly liquid markets—precisely the contexts where the relevance and applicability of such structurally rich models is typically more limited.
@@ 7 Conclusions
This work has presented a unified framework for understanding and optimizing dealer behavior in electronic bond trading on Multi-Dealer-to-Client (MD2C) platforms, grounded in the formal tools of probabilistic graphical models and causal inference. By representing the RFQ process through a causal graph, we uncover latent structures and confounding relationships that often undermine traditional data-driven approaches to business interventions such as pricing, client targeting, and profitability estimation. Our framework supports a rigorous analysis of these interventions, enabling the identification of conditions under which valid causal reasoning can be applied—leveraging the rich historical datasets typically available to dealers. In the absence of such tools, dealers risk deploying suboptimal strategies that fail to account for underlying causal dynamics.
@@
An empirical evaluation using a proprietary dataset of BBVA’s RFQ activity in European government bonds demonstrates that the generative model performs comparably to LightGBM in terms of AUC, while surpassing both discriminative models in calibration quality, as measured by the Balanced Brier Skill Score (BBSS). Although the logistic regression model offers interpretability and satisfies monotonicity by design, it underperforms in both classification and calibration metrics. This contrast is particularly noteworthy given that LightGBM, which is a state-of-the-art machine learning model, does not enforce monotonicity constraints by default.
@@
Several promising directions emerge for future work, offering pathways to extend and validate the proposed framework. One line of research focuses on enhancing the predictive performance of discriminative models by making full use of the available information, while also adhering to key business constraints. This includes developing models that incorporate domain-specific requirements—such as monotonic neural networks—and enhancing feature engineering to integrate informative post-trade signals like cover prices. This work is ongoing; see [26] for further developments.
Given the proprietary nature of the dataset used in this study, another important avenue is cross-institutional validation. Applying the framework to RFQ data from other institutions, even if non-public, would help assess its generality and robustness across different dealer environments and market structures.

### v2

@@ 6 Generative versus discriminative models for causal interventions: empirical results
We evaluate two discriminative models: a logistic regression and a tree-based model, the Light Gradient Boosting Machine (LightGBM). Logistic regression serves as a simple benchmark that inherently respects monotonicity constraints on the predicted probability with respect to input features—particularly the quoted spread. In contrast, LightGBM is a powerful, widely used model known for its strong out-of-the-box predictive performance. Although it does not enforce monotonicity, we employ it to explore the predictive frontier of the problem, assessing the potential gains in accuracy when the monotonicity constraint is relaxed. Both of them are efficient and interpretable, making them standard candidates to implement the hit probability model  [33]. A promising alternative involves neural network architectures that explicitly incorporate monotonicity constraints; see [31] for a related application in the context of derivatives pricing. These models aim to combine the principled structure of monotonic estimators with the flexibility and representational power of neural networks. Results from this line of research will be presented in future work [32].
@@
Finally, we include features specific to the RfQ level, corresponding to the variable $RF$ in the graphical model:
@@
We use a proprietary dataset of RfQs from BBVA, consisting of 102,437 RfQs, of which 5,738 are labeled as hits. More details about the data are available in S1 Table.. The dataset is split into training, validation, and test sets, with the test set reserved for final model evaluation. The validation set is used for model selection and hyperparameter tuning in the discriminative ML models. Full details of the preprocessing steps are provided in S1 Appendix.. For the logistic regression model, we tune the inverse regularization strength, setting it to 100 with an L2 penalty. For the LightGBM model, the best-performing hyperparameters identified are: column sample by tree = 1, learning rate = 0.01, minimum child samples = 50, number of estimators = 500, number of leaves = 15, and subsample = 0.6. The generative model does not require hyperparameter tuning and is trained using Maximum Likelihood Estimation (MLE) on the training set. Given the class imbalance in the dataset, we apply class weighting during training for both the generative and discriminative models. In addition, standard outlier cleaning techniques are applied to preprocess the dataset.
@@
A worst-case Balanced Brier Score (BBS) is $0.5$, while a perfect model achieves a score of 0. In terms of skill, the Balanced Brier Skill Score (BBSS) equals 0 when the model’s performance matches that of the benchmark, and reaches $1$ for a perfect model. This metric provides a succinct summary of the information conveyed by calibration plots, which group predicted probabilities into bins and compare the average predicted probability within each bin to the observed frequency of positive outcomes. Such plots—and, by extension, the Brier score—are particularly informative in settings where the predicted probabilities themselves, rather than just binary class labels, directly influence decision-making, as is the case in the business problems analyzed in this work.
The results for AUC-ROC and BBSS in the test set for the three models considered as well as the majority class benchmark are shown in Table 1.
Model AUC-ROC BBSS
@@
Table 1: Predictive scores: AUC-ROC and BBSS for the three models analyzed. We also show the metrics for the benchmark model, the majority class. Both the generative model and LightGBM achieve comparable predictive performance in terms of ROC-AUC, outperforming logistic regression. While the generative model exhibits lower calibration performance, as indicated by a reduced BBSS, it uniquely satisfies the key business constraint of spread monotonicity—an essential property for applications such as optimal pricing.
The comparative analysis of the three models reveals key insights into both predictive accuracy and the quality of probability estimates. LightGBM achieves the highest ROC-AUC at 0.743, but its advantage over the generative model (0.742) is marginal and not practically significant. Both models substantially outperform logistic regression, which attains a lower AUC of 0.684. These results are reinforced by inspecting the ROC curve in Fig 3.
In terms of calibration, however, the generative model exhibits a lower Balanced Brier Skill Score (BBSS) than LightGBM and Logistic Regression, indicating less accurate probability estimates. This trade-off is partly attributable to the model’s distributional assumptions, which constrain its flexibility compared to discriminative methods. But also visual inspection of the calibration plots shown in Fig 4 suggests that the generative model produces smooth, nearly monotone probability estimates. However, once we evaluate calibration with class-balanced weighting, the ranking reverses. This indicates that the generative model’s calibration errors concentrate on the minority (hit) class: it looks well-calibrated on average, but once errors are weighted equally across classes, logistic regression and LightGBM provide more accurate probabilities.
@@
Fig 4: Calibration plots for the three evaluated models and the majority class baseline. While the generative model appears smooth and close to the diagonal, class-balanced metrics indicate better calibration for logistic regression and LightGBM (see BBSS values in Table 1).
Despite this, the generative framework provides an important practical advantage: it inherently enforces the business-critical constraint of spread monotonicity, ensuring that predicted hit probabilities increase as quoted spreads become more competitive. This property, shared by construction with the Logistic Regression model, is essential for economically consistent applications such as optimal pricing and dealer quote recommendations. As illustrated in Fig 5, in contrast to the other two approaches, LightGBM exhibits erratic behavior in these areas, failing to reflect the expected increase in win probability as the quoted spreads become more attractive.
@@
Overall, these findings underscore the benefits of modeling approaches that capture the intrinsic structure and logic of the underlying economic process. Although purely predictive models may achieve incremental gains through advanced architectures, hyperparameter optimization, or feature enrichment with post-trade information, the observed performance ceiling suggests that predictive accuracy alone is insufficient. In contrast, models that incorporate economic reasoning and structural constraints can deliver more robust and interpretable outcomes without sacrificing predictive power. This advantage is particularly salient in complex financial settings such as MD2C negotiation platforms—systems that are sufficiently intricate to pose meaningful modeling challenges, yet structured enough to permit explicit representation of their internal mechanisms.
In the case of the generative model, this strength arises from its explicit representation of the internal RfQ negotiation mechanics—capturing, for example, the simultaneous bidding behavior of multiple dealers and the client’s selection of the most competitive quote. Furthermore, its ability to integrate post-trade signals, such as the cover price or information about being the second-best bid, enhances both interpretability and performance, reinforcing the value of structurally grounded modeling strategies in financial applications.
One potential limitation of structurally informed models, however, lies in their inference time. Generative approaches, in particular, tend to be slower than the other models evaluated. This drawback may be mitigated by the fact that low-latency inference is typically more critical in highly liquid markets—precisely the contexts in which the relevance and applicability of structurally rich models are often more limited. This consideration, though, helps explain why simpler approaches, such as logistic regression, remain popular in practical applications [1].
To better understand the drivers of predictive performance, we examine the importance of variables across models. In Table 2 we show the ranking of feature importance for each model. The majority class baseline does not use features, so it is not applicable. The ranking has been calculated using permutation importance for the generative model, standardized coefficients for the LR, and gain-based feature importance for LightGBM. More details are available in S3 Table..
Variable Generative Logistic R LightGBM
Spread 1 1 1
Volatility 3 7 8
DV01 5 2 7
Frequency buy 7 8 6
Frequency sell 6 6 9
Average dealers 4 4 5
Maturity 2 9 4
Number of dealers 8 5 3
DV01 exposure 9 3 2
Table 2: Ranking of feature importance across models for sell RfQs
Across models, spread is the dominant driver of hit probability: it ranks first in LightGBM, the logistic model, and the generative model. The second tier is led by risk/size proxies: DV01 (logistic) and DV01 exposure (LightGBM) carry substantial importance, indicating that larger risk‐adjusted notionals materially affect execution. Microstructure controls—number of dealers, maturity, volatility, and client frequency—play a secondary role: LightGBM assigns non-trivial importance to number of dealers (competition intensity) and to maturity (liquidity/duration risk), while the generative specification picks up maturity and volatility more than the frequency variables. The overall pattern is consistent: pricing aggressiveness (spread) and risk/size (DV01) are the primary drivers of execution, with platform and liquidity frictions shaping outcomes at the margin.
@@ 7 Conclusions
This work has presented a unified framework for understanding and optimizing dealer behavior in electronic bond trading on Multi-Dealer-to-Client (MD2C) platforms, grounded in the formal tools of probabilistic graphical models and causal inference. By representing the RfQ process through a causal graph, we uncover latent structures and confounding relationships that often undermine traditional data-driven approaches to business interventions such as pricing, client targeting, and profitability estimation. Our framework supports a rigorous analysis of these interventions, enabling the identification of conditions under which valid causal reasoning can be applied—leveraging the rich historical datasets typically available to dealers. In the absence of such tools, dealers risk deploying suboptimal strategies that fail to account for underlying causal dynamics. A summary of a typical workflow using our framework is provided in Fig 6.
[…]
[width=0.5]Fig6
Fig 6: Overview of the frameworks developed in this work for causal interventions on MD2C platforms, exemplified through the optimal pricing problem. We apply causal inference to identify the conditioning set $\mathcal{Z}_{t}$ that enables reduction of the problem to empirically estimable probabilities—specifically, the hit probability model. Both generative and discriminative model formulations are used to estimate this probability, with parameters fitted via maximum likelihood. Their performance is then evaluated on a separate validation set using standard predictive metrics tailored to this problem, while incorporating business constraints such as monotonicity with respect to the spread. Finally, the selected model is employed to compute the optimal spread, denoted $\delta^{*}$
@@
An empirical evaluation based on a proprietary dataset of BBVA’s Request-for-Quote (RfQ) activity in European government bonds shows that the generative model achieves performance comparable to LightGBM in terms of AUC. The logistic regression model, while offering straightforward interpretability and satisfying monotonicity by construction, underperforms on classification metrics. This contrast is particularly noteworthy given that LightGBM—a state-of-the-art machine learning approach—does not impose monotonicity constraints by default. The trade-off for the generative model’s structural rigor is a deterioration in calibration metrics, as its reduced flexibility limits its ability to produce well-calibrated probabilities across the entire range. Nevertheless, for the dealer intervention use cases discussed in this study, we consider this trade-off worthwhile, since violating structural constraints such as monotonicity can lead to economically inconsistent or unreliable decisions in practice.
@@
Several promising directions emerge for future work, offering pathways to extend and validate the proposed framework. One line of research focuses on enhancing the predictive performance of discriminative models by making full use of the available information, while also adhering to key business constraints. This includes developing models that incorporate domain-specific requirements—such as monotonic neural networks—and enhancing feature engineering to integrate informative post-trade signals like cover prices. This work is ongoing; see [32] for further developments. Other recent studies have explored alternative feature engineering approaches, including quantum-enhanced representations that aim to enrich probabilistic models with higher-order informational structure [35].
Given the proprietary nature of the dataset used in this study, an important direction for future work is cross-institutional validation. Applying the proposed framework to RfQ data from other institutions—even if non-public—would help evaluate its generality and robustness across different dealer environments and market structures. Our results indicate that the causal intervention framework enhances predictive modeling in bond RfQ platforms. When the trading mechanism remains the same—i.e., an RfQ-based dealer-to-client process—the framework should be overall portable; what requires adaptation is the conditioning set and, potentially, the parameterization. Extending the approach to other asset classes, such as corporate bonds or derivatives, will therefore necessitate re-tuning the instrument characteristics, and client profiles. For example, the conditioning variables may need to incorporate credit ratings, alternative risk measures (e.g., CDS spreads, duration/DV01), collateral eligibility, margining practices, or contract-specific features (e.g., option Greeks). In some cases, minor variations in the RfQ trading protocol may need to be adapted to the specific use case. In summary, while the graphical model is, in principle, flexible enough to accommodate most of these elements, a careful reassessment would still be required before applying the framework to new contexts.
@@
Supporting information
S1 Appendix.
Data preprocessing. Complete description of cleaning, filters, and outlier removal.
S1 Table.
Data summary table. Summary statistics of the RFQ dataset (size, period, sides, dealers, notional, frequency).
S2 Table.
Notation. Reference table.
S3 Table.
Feature importance. Feature importance across models for sell RfQs.